# Strong Data Processing Properties of Rényi-divergences via Pinsker-type Inequalities

### v2

@@ Appendix A Proof of Proposition 3
$\sup_{P_{X},Q_{X}\in\mathcal{P}_{X}:\,D_{\alpha}(P_{X}||Q_{X})\neq 0}\frac{\log\biggl{\{}\sum_{y\in\mathcal{Y}}\biggl{(}\sum_{x\in\mathcal{X}^{(y)}}P_{Y|X=x}(y)P_{X}(x)\biggr{)}^{\alpha}\biggl{(}\sum_{x\in\mathcal{X}^{(y)}}P_{Y|X=x}(y)Q_{X}(x)\biggr{)}^{1-\alpha}\biggr{\}}}{\log\biggl{\{}\sum_{x\in\mathcal{X}}P_{X}(x)^{\alpha}Q_{X}(x)^{1-\alpha}\biggr{\}}}$ (49)
@@
$\frac{\log\biggl{\{}\sum_{y\in\mathcal{Y}}\biggl{(}\sum_{x\in\mathcal{X}^{(y)}}P_{Y|X=x}(y)P_{X}(x)\biggr{)}^{\alpha}\biggl{(}\sum_{x\in\mathcal{X}^{(y)}}P_{Y|X=x}(y)Q_{X}(x)\biggr{)}^{1-\alpha}\biggr{\}}}{\log\biggl{\{}\sum_{x\in\mathcal{X}}P_{X}(x)^{\alpha}Q_{X}(x)^{1-\alpha}\biggr{\}}}=1.$ (50)
Since the DPI for Rényi-divergences states that $\eta_{\alpha}(P_{Y|X})\leq 1$ for all $P_{Y|X}$, constructing this distribution proves $\eta_{\alpha}(P_{Y|X})=1$.
@@
where $j$ is such that any $\mathcal{X}^{(y)}\coloneqq\{x\in[N]:P_{Y|X=x}(y)>0\}$ that contains $i$ does not contain $j$. Then we get
$\lim_{\gamma\to 0}\eta_{\alpha}(P^{\gamma}_{X},P_{Y|X})$ $=\frac{\log\biggl{\{}\sum_{y\in\textnormal{supp}(P_{Y|X=i})}\biggl{(}P_{Y|X=i}(y)Q^{0}_{X}(i)\biggr{)}^{1-\alpha}\biggl{(}P_{Y|X=i}(y)\biggr{)}^{\alpha}\biggr{\}}}{\log\biggl{\{}Q_{X}^{0}(i)^{1-\alpha}\biggr{\}}}$ (54)
$=\frac{\log\biggl{\{}Q^{0}_{X}(i)^{1-\alpha}\sum_{y\in\textnormal{supp}(P_{Y|X=i})}P_{Y|X=i}(y)\biggr{\}}}{\log\biggl{\{}Q^{0}_{X}(i)^{1-\alpha}\biggr{\}}}$ (55)
@@
where the second to last equality follows from the row-stochasticity of $P_{Y|X}$.
By this construction we have $\eta_{\alpha}(P_{Y|X})\geq 1$. Hence, the data processing inequality for Rényi-divergences implies $\eta_{\alpha}(P_{Y|X})=1$. Noticing that choosing $i$ and $j$ in this manner is only possible if there exists some $i,j\in[N]$ such that $\textnormal{supp}(P_{Y|X=i})\cap\textnormal{supp}(P_{Y|X=j})=\emptyset$ finishes the proof. ∎
Appendix B Proof of Theorem 8
where $d_{f_{\alpha}}(p||q)$ denotes the binary $f_{\alpha}$-divergence
From the definition of the total variation, we get
$\min_{p,q:\,|p-q|=t}d_{f_{\alpha}}(p||q)=\min_{p\in[0,1-t]}\biggl{\{}\frac{p^{\alpha}}{(p+t)^{\alpha-1}}+\frac{(1-p)^{\alpha}}{(1-p-t)^{\alpha-1}}\biggr{\}}-1.$ (59)
Define
Taking derivatives with respect to $p$ yields
and for any $\alpha>1$ and all $t<1,\,p\leq 1-t$,
from which we can conclude that $\psi_{\alpha}(p,t)$ is convex in $p$.
with equality for $t=\nicefrac{{1}}{{\alpha}}$ if $p=0$. Otherwise strict inequality holds. Together with the convexity of $\psi_{\alpha}$ in $p$, this implies that the maximum value of $\psi_{\alpha}$ for all $t\geq\nicefrac{{1}}{{\alpha}}$ will be achieved at $p=0$. At this point, we have
$\psi_{\alpha}(0,t)=(1-t)^{1-\alpha}.$ (64)
$=\biggl{(}1+\frac{t^{2}}{(p+t)(1-p-t)}\biggr{)}^{\alpha-1}-1$
$\geq(1+4t^{2})^{\alpha-1}-1.$
Finally, we bound the remaining case with the standard Pinsker’s inequality: Since the Rényi-divergence is increasing in $\alpha$ [9], so is $D_{f_{\alpha}}$. Hence the standard Pinsker’s inequlity for the Kullback-Leibler divergence $\text{KL}(P||Q)$ (Rényi-divergence of order $\alpha=1$)
$\text{KL}(P||Q)=D_{1}(P||Q)\geq 2\text{TV}(P||Q)^{2}$ (65)
will continue to hold for $\alpha>1$. Utilizing (4), this yields
Note a tight bound in the case $t<\nicefrac{{1}}{{\alpha}}$ can be obtained by solving
$\frac{\partial}{\partial p}\psi_{\alpha}(p,t)=(\alpha t+p)\biggl{[}\frac{p^{\alpha-1}}{(p+t)^{\alpha}}+\frac{(1-p)^{\alpha-1}}{(1-p-t)^{\alpha}}\biggr{]}-\frac{(1-p)^{\alpha-1}}{(1-p-t)^{\alpha}}=0.$ (67)
In the range $t\in(0,\nicefrac{{1}}{{\alpha}})$, this equation has a real solution $p^{*}$ that satisfies $p^{*}\in[0,1]$. Due to the convexity of $\psi_{\alpha}$ in $p$, a (tight) lower bound is then given by
$g_{\alpha}(t)=\psi_{\alpha}(p^{*},t).$ (68)
However, it remains unclear if $p^{*}$ can be expressed in closed form. Due to this complication, the result above utilizes more rudimentary bounds in this case.
Appendix C Proof of Proposition 12
Note that for a randomized response mechanism according to [31], we can write for any two rows $w$, $w^{\prime}$,
$P_{X|W=w}(x)=P_{X|W=w^{\prime}}(x)+\kappa(x),$ (69)
where
$\kappa(x)\in\{0,\frac{e^{\varepsilon}-1}{|\mathcal{X}|+e^{\varepsilon}-1},\frac{1-e^{\varepsilon}}{|\mathcal{X}|+e^{\varepsilon}-1}\}$ (70)
and each non-zero element is taken exactly once. Therefore,
$||\kappa||_{1}=2\frac{e^{\varepsilon}-1}{|\mathcal{X}|+e^{\varepsilon}-1}.$ (71)
Using this observation, we can write
$\Gamma_{\max}$ $=\max_{w,w^{\prime}}\,\max_{y}\frac{\sum_{x}P_{Y|X=x}(y)P_{X|W=w^{\prime}}(x)}{\sum_{x}P_{Y|X=x}(y)P_{X|W=w}(x)}$ (72)
$=\max_{\kappa:\,||\kappa||_{1}=2\frac{e^{\varepsilon}-1}{|\mathcal{X}|+e^{\varepsilon}-1}}\max_{w^{\prime}}\,\max_{y}\frac{\sum_{x}P_{Y|X=x}(y)P_{X|W=w^{\prime}}(x)}{\sum_{x}P_{Y|X=x}(y)P_{X|W=w^{\prime}}(x)+\sum_{x}P_{Y|X=x}(y)\kappa(x)}$ (73)
$=\max_{w^{\prime}}\,\max_{y}\frac{\sum_{x}P_{Y|X=x}(y)P_{X|W=w^{\prime}}(x)}{\sum_{x}P_{Y|X=x}(y)P_{X|W=w^{\prime}}(x)-\frac{||\kappa||_{1}}{2}\biggl{(}\max_{x}P_{Y|X=x}(y)-\min_{x}P_{Y|X=x}(y)\biggr{)}}$ (74)
$\leq\max_{w^{\prime}}\,\max_{y}\frac{\sum_{x}P_{Y|X=x}(y)P_{X|W=w^{\prime}}(x)}{\sum_{x}P_{Y|X=x}(y)P_{X|W=w^{\prime}}(x)-\frac{||\kappa||_{1}}{2}\max_{x}P_{Y|X=x}(y)}.$ (75)
Now, fix $y$ and let $w^{*}\in\mathcal{W}$ be the solution to the maximization over $\mathcal{W}$. By letting
$\ell_{w^{*}}(X\to y)=\log\frac{\max_{x}P_{Y|X=x}(y)}{\sum_{x}P_{Y|X=x}(y)P_{X|W=w^{*}}(x)},$ (76)
we can further simplify the above bound as
$\Gamma_{\max}\leq\max_{y}\frac{1}{1-\frac{||\kappa||_{1}}{2}\exp\ell_{w^{*}}(X\to y)}=\biggl{(}1-\frac{e^{\varepsilon}-1}{|\mathcal{X}|+e^{\varepsilon}+1}\exp{\max_{y\in\mathcal{Y}}\min_{w\in\mathcal{W}}}\,\ell_{w}(X\to y)\biggr{)}^{-1}.$ (77)
The proof for $\Gamma_{\min}$ follows analogously. ∎

### v3

@@ Appendix B Proof of Proposition 3
$\sup_{P_{X},Q_{X}\in\mathcal{P}_{X}:\,D_{\alpha}(P_{X}||Q_{X})\neq 0}\frac{\log\bigg{\{}\sum_{y\in\mathcal{Y}}P_{Y}(y)^{\alpha}Q_{Y}(y)^{1-\alpha}\biggr{\}}}{\log\biggl{\{}\sum_{x\in\mathcal{X}}P_{X}(x)^{\alpha}Q_{X}(x)^{1-\alpha}\biggr{\}}}$ (71)
@@
$\frac{\log\biggl{\{}\sum_{y\in\mathcal{Y}}P_{Y}(y)^{\alpha}Q_{Y}(y)^{1-\alpha}\biggr{\}}}{\log\biggl{\{}\sum_{x\in\mathcal{X}}P_{X}(x)^{\alpha}Q_{X}(x)^{1-\alpha}\biggr{\}}}=1.$ (72)
Since the DPI for Rényi-divergences states that $\eta_{\alpha}(P_{Y|X})\leq 1$ for all $P_{Y|X}$, this then implies $\eta_{\alpha}(P_{Y|X})=1$.
@@
where $j$ is such that any $\mathcal{X}^{(y)}\coloneqq\{x\in[N]:P_{Y|X=x}(y)>0\}$ that contains $i$ does not contain $j$. Let $\mathcal{S}_{i}\coloneqq\textnormal{supp}(P_{Y|X=i})$. Then we get
$\lim_{\gamma\to 0}\eta_{\alpha}(P^{\gamma}_{X},P_{Y|X})$ (76)
$=\frac{\log\biggl{\{}\sum_{y\in\mathcal{S}_{i}}\biggl{(}P_{Y|X=i}(y)Q^{0}_{X}(i)\biggr{)}^{1-\alpha}\biggl{(}P_{Y|X=i}(y)\biggr{)}^{\alpha}\biggr{\}}}{\log\biggl{\{}Q_{X}^{0}(i)^{1-\alpha}\biggr{\}}}$ (77)
$=\frac{\log\biggl{\{}Q^{0}_{X}(i)^{1-\alpha}\sum_{y\in\mathcal{S}_{i}}P_{Y|X=i}(y)\biggr{\}}}{\log\biggl{\{}Q^{0}_{X}(i)^{1-\alpha}\biggr{\}}}$ (78)
@@
where the second to last equality follows from the row-stochasticity of $P_{Y|X}$. By this construction we have $\eta_{\alpha}(P_{Y|X})\geq 1$. Hence, the data processing inequality for Rényi-divergences implies $\eta_{\alpha}(P_{Y|X})=1$. Noticing that choosing $i$ and $j$ in this manner is only possible if there exists some $i,j\in[N]$ such that $\textnormal{supp}(P_{Y|X=i})\cap\textnormal{supp}(P_{Y|X=j})=\emptyset$ finishes the proof. ∎